# Multi-Objective Bayesian Optimization for Networked Black-Box Systems: A Path to Greener Profits and Smarter Designs

### v1

@@ 1 Introduction
Recent work in so-called “composite” BO has sought to extend BO to systems where one function’s output feeds into another [35, 26, 29]. However, these methods assume that the function can be expressed as a composite model, $g(\bm{h}(\bm{x}))$, where $g$ is known, $\bm{h}$ is a black-box function, and $\bm{x}$ are the design variables. A more general representation of the function would be as a network (or graph), which has been explored in recent BO frameworks such as BOFN [36] and BOIS [37, 38]. Graph-based models are known to be powerful representations are seen by their wide applicability in many different fields such as social sciences, biology, chemistry, and engineering [39]. However, these current graph-based BO methods assume that the function network is a directed acyclic graph (DAG) to simplify the uncertainty propagation task. This assumption limits their applicability to process systems with feedback loops, such as recycle streams and interconnected multi-scale simulations.
@@
To ensure computational efficiency of the algorithm, MOBONS extends the Thompson sampling acquisition function [40], which is significantly cheaper than the improvement-based approach in BOFN [36] and the confidence bound approach in BOIS [37, 38]. Furthermore, we leverage multi-objective genetic algorithms to efficiently solve the subproblems within MOBONS, combining the sample efficiency of BO with the exploration power of genetic search.
@@ 2 Problem Formulation
[…]
@@ 3.1 Statistical surrogate model
where $\mu_{0,k}(\bm{z}_{k}):\mathbb{R}^{n_{z,k}}\to\mathbb{R}$ is the prior mean function and $\Sigma_{0,k}(\bm{z}_{k},\bm{z}_{k}^{\prime}):\mathbb{R}^{n_{z,k}}\times\mathbb{R}^{n_{z,k}}\to\mathbb{R}_{+}$ is the prior covariance (or kernel) function (note $n_{z,k}$ denotes the dimension of $\bm{z}_{k}$ and $\mathbb{R}_{+}$ denotes the set of positive real numbers). The mean function encodes prior knowledge about the expected behavior of $f_{k}$, while the kernel function defines correlations between function values at different inputs, capturing smoothness and variability. A detailed review of kernel choices can be found in [47, Chapter 4]; in this work, we assume an appropriate kernel is selected, potentially learned from data.
@@ 3.2 Multi-objective Thompson sampling for function networks
[…]
@@
[…]
@@ 3.3 Practical considerations in MOBONS
To formulate the acquisition function (10), we require posterior function samples from the GP models. The complexity of this step depends on the cost of evaluating $\widehat{\bm{F}}_{t}$. To generate consistent (differentiable) posterior samples, we recommend the efficient procedure in [50], which decomposes a GP posterior sample into two steps: (i) Drawing a realization from the weight-space version of the prior and (ii) Performing a function-space update based on observed data. This requires knowledge of the weight-space interpretation of the GP, where the covariance function is viewed as an inner product in a feature space. For stationary kernels, this feature space can be approximated using random Fourier features (RFF) [51], enabling scalable sampling. A detailed discussion of RFF-based posterior sampling is provided in [45, Section 3].
@@ 4.2 Synthetic test problem: ZDT4 benchmark
The feasible set for the design variables is constrained within the hyperrectangle: $0\leq x_{1}\leq 1$ and $-10\leq x_{k}\leq 10$ for $k=2,\ldots,10$. We assume that the node functions $f_{1,\dots,10}$ are black-box functions (expensive and unknown), while $f_{11}$ is a white-box function (computationally inexpensive and known).
@@
where $[\bm{r},\bm{p}_{i}]$ is the hyper-rectangle defined by the reference point $\bm{r}$ and a Pareto-optimal solution $\bm{p}_{i}$. For this case study, we select a reference point of $\bm{r}=[1,500]$. Each algorithm is initialized with $n=21$ randomly chosen samples in $\mathcal{X}$, and the total evaluation budget is $N=121$ (100 additional sequentially selected points). To account for stochastic variability due to random initialization, we repeat all experiments 30 times, ensuring consistent random seeds across all algorithms.
@@ 4.3 Design of sustainable bioethanol process
Another advantage of MOBONS is that it produces a surrogate model that incorporates network structure, enabling additional analyses such as sensitivity analysis (SA). Here, we use SA to examine the impact of different decision variables on the two objectives (revenue and GWP). Since the function network GP model may not be globally accurate, we focus on a local SA around solutions along the Pareto frontier. We conduct local SA on four selected Pareto solutions (points A, B, C, and D in Figure 8) by perturbing each decision variable within a ±$\pm 5$% range. To improve model accuracy in these local regions, we generate an additional 24 function evaluations centered around each Pareto point. First-order Sobol indices [58] are then computed to quantify the relative influence of each decision variable on revenue and GWP. Figure 9 presents the results of the local SA. We see that revenue is primarily influenced by water/glucose feed flowrate ($F$) and ethanol purity in the first distillation column ($p$). Notably, as revenue increases along the Pareto front (moving from left to right), its sensitivity to $F$ grows, while its sensitivity to $p$ diminishes. On the other hand, we see that GWP is predominantly impacted by $F$, with negligible (local) sensitivity to other variables. This result aligns with expectations, as CO2 emissions scale directly with the amount of material fed into the process.

### v2

@@ 1 Introduction
Recent work in so-called “composite” BO has sought to extend BO to systems where one function’s output feeds into another [35, 26, 29]. However, these methods assume that the function can be expressed as a composite model, $g(\bm{h}(\bm{x}))$, where $g$ is known, $\bm{h}$ is a black-box function, and $\bm{x}$ are the design variables. A more general representation of the function would be as a network (or graph), which has been explored in recent BO frameworks such as BOFN [36, 37] and BOIS [38, 39]. Graph-based models are known to be powerful representations are seen by their wide applicability in many different fields such as social sciences, biology, chemistry, and engineering [40]. However, these current graph-based BO methods assume that the function network is a directed acyclic graph (DAG) to simplify the uncertainty propagation task. This assumption limits their applicability to process systems with feedback loops, such as recycle streams and interconnected multi-scale simulations.
@@
To ensure computational efficiency of the algorithm, MOBONS extends the Thompson sampling acquisition function [41], which is significantly cheaper than the improvement-based approach in BOFN [36] and the confidence bound approach in BOIS [38, 39]. Furthermore, we leverage multi-objective genetic algorithms to efficiently solve the subproblems within MOBONS, combining the sample efficiency of BO with the exploration power of genetic search. Note that we have also recently pursued related work in the context of function networks under adversarial uncertainty (i.e., robust optimization) [42], though we do not directly consider that case in this work.
@@ 2 Problem Formulation
Building on recent work on function networks [36], we consider a structured representation of $\bm{G}$ in which the objectives are evaluated through a directed network of interdependent functions $f_{1},\ldots,f_{K}$. Each function $f_{k}$ in the network receives inputs from a subset of external design variables $\bm{x}$ as well as the outputs of other functions in the network. The evaluation of these functions is assumed to be expensive, meaning that computational or experimental resources are required for function evaluations. To formalize this structure, let $V=\{1,2,\ldots,K\}$ represent the set of nodes in the network, where each node corresponds to a function $f_{k}$. The directed edges of the network, denoted by $E=\{(j,k):f_{k}\text{~receives input from~}f_{j}\}$, specify the dependencies between functions. Each function $f_{k}$ can be expressed as an implicit equation
@@ 3.1 Statistical surrogate model
where $\mu_{0,k}(\bm{z}_{k}):\mathbb{R}^{n_{z,k}}\to\mathbb{R}$ is the prior mean function and $\Sigma_{0,k}(\bm{z}_{k},\bm{z}_{k}^{\prime}):\mathbb{R}^{n_{z,k}}\times\mathbb{R}^{n_{z,k}}\to\mathbb{R}_{+}$ is the prior covariance (or kernel) function (note $n_{z,k}$ denotes the dimension of $\bm{z}_{k}$ and $\mathbb{R}_{+}$ denotes the set of positive real numbers). The mean function encodes prior knowledge about the expected behavior of $f_{k}$, while the kernel function defines correlations between function values at different inputs, capturing smoothness and variability. A detailed review of kernel choices can be found in [49, Chapter 4]; in this work, we assume an appropriate kernel is selected, potentially learned from data.
@@ 3.2 Multi-objective Thompson sampling for function networks
$p_{\bm{x}^{\star}}(\bm{x}|\mathcal{D}_{t})=\int\delta\left(\bm{x}-\operatorname*{argmin}_{\bm{x}\in\mathcal{X}}g(\bm{x})\right)p(g|\mathcal{D}_{t})dg,$ (7)
@@
$=\int\mathbb{I}\left((\bm{x},\bm{Y})\in\operatorname*{argmin}_{\bm{x}\in\mathcal{X},\bm{Y}}\{\bm{C}\bm{Y}:\bm{Y}=\bm{F}(\bm{x},\bm{Y})\}\right)p(f_{1}|\mathcal{D}_{t,1})\cdots p(f_{K}|\mathcal{D}_{t,K})df_{1}\cdots df_{K},$
@@ 3.3 Practical considerations in MOBONS
To formulate the acquisition function (10), we require posterior function samples from the GP models. The complexity of this step depends on the cost of evaluating $\widehat{\bm{F}}_{t}$. To generate consistent (differentiable) posterior samples, we recommend the efficient procedure in [52], which decomposes a GP posterior sample into two steps: (i) Drawing a realization from the weight-space version of the prior and (ii) Performing a function-space update based on observed data. This requires knowledge of the weight-space interpretation of the GP, where the covariance function is viewed as an inner product in a feature space. For stationary kernels, this feature space can be approximated using random Fourier features (RFF) [53], enabling scalable sampling. A detailed discussion of RFF-based posterior sampling is provided in [47, Section 3].
@@ 4.2 Synthetic test problem: ZDT4 benchmark
The feasible set for the design variables is constrained within the hyperrectangle: $0\leq x_{1}\leq 1$ and $10\leq x_{k}\leq 10$ for $k=2,\ldots,10$. We assume that the node functions $f_{1,\dots,10}$ are black-box functions (expensive and unknown), while $f_{11}$ is a white-box function (computationally inexpensive and known).
@@
where $\bm{r},\bm{p}_{i}]$ is the hyper-rectangle defined by the reference point $\bm{r}$ and a Pareto-optimal solution $\bm{p}_{i}$. For this case study, we select a reference point of $\bm{r}=[1,500]$. Each algorithm is initialized with $n=21$ randomly chosen samples in $\mathcal{X}$, and the total evaluation budget is $N=121$ (100 additional sequentially selected points). To account for stochastic variability due to random initialization, we repeat all experiments 30 times, ensuring consistent random seeds across all algorithms.
@@ 4.3 Design of sustainable bioethanol process
Another advantage of MOBONS is that it produces a surrogate model that incorporates network structure, enabling additional analyses such as sensitivity analysis (SA). Here, we use SA to examine the impact of different decision variables on the two objectives (revenue and GWP). Since the function network GP model may not be globally accurate, we focus on a local SA around solutions along the Pareto frontier. We conduct local SA on four selected Pareto solutions (points A, B, C, and D in Figure 8) by perturbing each decision variable within a $\pm 5$% range. To improve model accuracy in these local regions, we generate an additional 24 function evaluations centered around each Pareto point. First-order Sobol indices [60] are then computed to quantify the relative influence of each decision variable on revenue and GWP. Figure 9 presents the results of the local SA. We see that revenue is primarily influenced by water/glucose feed flowrate ($F$) and ethanol purity in the first distillation column ($p$). Notably, as revenue increases along the Pareto front (moving from left to right), its sensitivity to $F$ grows, while its sensitivity to $p$ diminishes. On the other hand, we see that GWP is predominantly impacted by $F$, with negligible (local) sensitivity to other variables. This result aligns with expectations, as CO2 emissions scale directly with the amount of material fed into the process.
@@ 5 Conclusion
Acknowledgments
The authors gratefully acknowledge the support from the Ohio Department of Higher Education Harmful Algal Bloom Research Initiative (ODHE HABRI).